40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 130 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  10424   Fri Aug 22 15:11:55 2014 andres, nicolasSummaryIOOMC WFS activity

1. Before doing anything, we centered the IOO QPDs.
2. With the WFS enabled, we offloaded the control signals onto the bias sliders. Then we saved the slider values. The MC LSC diode had a DC value of ~0.5
3. Turned down power with half wave plate before PMC.  Power injected to vacuum ~ 100mW.
4. We did a beam scan of MC REFL, it looks smaller than what Andres predicted based on the MC eigenmode by 10-20%.
5. We made many changes on the table, pictures to be added by Andres.
6. We didn't have the 80% reflector we wanted to increase the WFS power, so it's still a 98%.
6. Beams were aligned on MC REFL PL, camera, beam dumps, WFSs.
7. Clean up
8. PSL power increased to 1.2W, MC locked right away.
9 We didn't change the IOO WFS output matrix, but we changed some signs and gains to make everything stable. MC autolocker brings it back from cold just fine.
10. All time bombs that we've left will be E.Q.'s to clean up. Sorry.\
11. Yay

  10423   Fri Aug 22 13:51:00 2014 JenneUpdateGeneralUpdated game plan

40mToDoList.pdf

  10422   Fri Aug 22 03:55:45 2014 JenneUpdateLSCXarm PDH fine, Yarm PDH/ALS needs work

[Rana, Jenne, EricQ]

We did several things tonight.  First, a list (so I can remember them all), and then some details.

(1) Jiggled ETMY SUS cables, removed kicks.

(2) Locked X and Y ALS, looked at POX, POY as out of loop sensors.

(3) Measured stuff (?) at the Yend.

(4) Reconnected REFL DC to SR560.

(5) Attempted CARM offset reduction.


Item 1:

When Rana and I started locking this evening, we saw (as Q has been witnessing for a while now) the ETMY kick a lot.  However, it seemed to be kicking even more than usual.  Since Q had been down at the end station recabling things, we wondered if a SUS-related cable got bumped.  Rana went down to the end and pushed all the cables into their receptacles.  One of the last sets that he pushed was the satellite box.  We didn't have walkie-talkie communication, but the DC offset of the ETMY oplevs changed just a minute or two before he returned to the control room.  So, we guess that it was the satellite box cables that were loose.  Unfortunately, there is no clear way to strain relieve them, which is why they can so often be troublesome.  Anyhow, the ETMY hasn't kicked since.

Item 2:

We locked the arms with ALS.  We saw that the POX signal was about 20% of the full pk-pk height of the PDH signal, so it's mostly within the linear range, but not entirely.  It is what it is, however, and we took measurements assuming that it's okay.  I calibrated POX by putting an excitation onto ETMX, and matching the height of the peak in POX and BEATX_FINE_PHASE_OUT_HZ.

Q and Rana had also [remembered / put in / something] a digital readback for the end green PDH error point.  Q went down to the end and gave me a number of 2600 Hz/V for the err mon port of the PDH board, which is what is connected to the ADC.  With that and 20/2^16 V/cts, I had a calibration of 0.8 Hz/ct. 

What we see in this plot is that the green end PDH is not the limiting noise for the POX out of loop measurement of the residual arm motion.  Also, in the multi-color metrology paper, Fig 7 (which is posted in the control room), we see at about a little over 1 Hz a ratio of about 4.5 between the residual motion and the AUX PDH error signal.  In today's plot, I see a ratio of about 20.  I infer from this that the green PDH for the Xarm is fine, and that we may want to re-look at the ALS digital loop, but we should leave the X PDH alone.

Here is the Xarm plot:

Xend_ErrorPointMeasurements.pdf

Q took the data for the Yarm plot, so hopefully he can give it to us in the morning.  What we did notice was that the noise was much worse for the Yarm.  This prompted Item 3, measuring the loop.

Item 3:

Q and Rana went down to the Yend and measured some things.  They came back, and said that they hadn't changed anything in analog while they were down there.  One thing that Q did note was that we have almost 90 degrees of phase margin (since it's a 1/f loop), and about 10 dB of gain margin, above the UGF.  So, we're in good shape for being able to try triggering the boost on the PDH box.  Q will give us more notes on this work, as well as plots, in the morning.

Item 4:

At some point, I remembered that Q and Gabriele had repurposed the SR560 that we had been using for the REFLDC input to the common mode board.  So, Q went and put it back, so that REFL DC goes into the SR560, and so does a DAC channel so that we can remotely set the offset.  The A-B output goes to the REFL11I whitening channel, since real REFL11I goes into the input of the CM board.  I think that today, the SR 560 was left at a gain of 1.

Item 5:

We decided to carry on and try to reduce the CARM offset some.  An annoyance is that the Yarm still has pretty significant low-frequency noise, but the idea is that if we can get over to the sqrtInvTrans signals, it will be fine.

So, we didn't get much farther than we had in the past, but it was nice to get there at all again.  I ran the carm_cm_up script (many times).  One of the times, all I wanted to do was see how much I could reduce the CARM offset.  CARM was on sqrtInvTrans, DARM was on ALS diff, and I was able to get the arm powers up to about 2.5.  I don't know why I lost lock.  The sqrtInv signals should be good until at least arm powers of 20 or so. 

I was able to see the REFL DC dip, but only a teensy tiny bit.  It went down by maybe 1 count.  Q suggested looking at how deep it could get while leaving CARM and DARM both on ALS, and setting both offsets to 0.  We were seeing arm flashes of about 50 counts, and REFL DC went from 0 to -800.  So, I wasn't seeing much of a REFL dip, but it was definitely there when I went to arm powers of 2ish.

We tried looking at different sqrtInv options for DARM, and haven't come to any real conclusion.  In the plot below, we are looking at a swept sine between DARM_IN1 (ALSdiff) and either MC_IN1 0.3*(sqrtInvX - sqrtInvY) or SRCL_IN1 (TRX - TRY / sqrt(TRX + TRY) ):

DARM_ALSdiff_vs_sqrtInv.pdf

 


We have a few things to add to the to-do list:

* Put UGF servos for LSC loops in place.

* Implement UGF "servos" (per Koji's suggested method) for phase trackers.

* Write a lockloss script that is run by the ALS watch scripts - print a PDF of error and control signals for every lockloss, and save it somewhere.

* Fix up Ygreen modematching on the PSL table.  The X green spot is quite similar on the camera to the corresponding PSL green spot.  However the Y green spot is not at all the same as its PSL green spot. 

 

  10421   Thu Aug 21 22:10:52 2014 ranaSummaryComputer Scripts / Programsnetwork movements

 To help development of the data visualization project, we've assigned the .101 and .102 IP to DataVis. This is being used by the iMac in the control room via port 8 of the CDS switch near the Blue Plataeu Tournant.

We tried using one of the free ports, but Jamie realized that we had to use one of the already assigned ones due to some 'Smart' switch management software. So for the moment, please leave the iMac alone so that Bill can use it.

  10420   Thu Aug 21 19:04:52 2014 ericqConfigurationGreen LockingGain changes on Green Y PDH

I found that the barrel of one the BNC to BNC connectors used for getting the output of the PDH servo box to the laser controller was touching the ETMY chamber. When I held it away, all of the 60Hz harmonics disappeared from the mixer output spectrum; this was pretty repeatable. This inspired me to replace the refl PD and PZT signal cables (which were 2 and 3 cables stitched together, respectively) with 20' long BNCs. I also cleaned up a lot of the routing of signal and power cables in the little rack, and moved the big T->DC Block->Attenuator combo off of the panel mount, because I didn't like how it was wiggling. It and the summing pomona box are sitting on top of the PDH box and function generator, instead of hanging freely.

All of the 60Hz harmonics were banished afterwards, and the green locked happily. 

This required me touching the Y end table, to remove the old cable and its cable ties, and putting the new one in. I don't think I did anything immediately apparently bad; the green and IR transmissions both are within nominal ranges. 

I haven't had luck measuring the CLG yet, which I wanted to do to get and set the UGF before measuring the noises. However, here is a scope trace of the in-lock error signal, which compares quite favorably to the trace posted in the previous post; the scope indicates that the signal has 1/3 of the RMS that it did before I replaced the cables. 

TEK00005.PNG

I hope to measure up the current status after I get back from dinner. 

 

  10419   Thu Aug 21 15:07:48 2014 SteveUpdateVACRGA scan at day 197

 

 

Attachment 1: RGAscan@d197.png
RGAscan@d197.png
  10418   Thu Aug 21 02:42:17 2014 rana, ericqConfigurationGreen LockingGain changes on Green Y PDH

[rana, ericq]

We spent time trying to relieve the Yend green PDH of it troubles. 

We realized that the mixer in the PDH setup (mini circuits ZAD-8+), wants 7dBm of LO to properly function. However, we use one function generators output, through a splitter, to give signals to the laser PZT and the mixer LO. 

We don't want 7dBm of power hitting the laser PZT, though. The summing node that adds the servo output to the sideband signal was supposedly designed to do some of this attenuation. Rana measured that 10Vpp out of the function generator resulted in 20mVpp on the fast input to the NPRO, after the summing node. Hence, the 0.09V setting was only resulting in something like 0.2mV hitting the PZT. The PZT has something like 30 rad/V PM response, meaning we only had ~0.006 rad of modulation. 

Now, the function generator is set to 2 Vpp, meaning 4 mVpp hitting the PZT, meaning ~0.12 radians of modulation. The mixer is now getting +7dBm on its LO, and the PDH traces look much cleaner. However, the PDH error signal is now something like 100mVpp, which is much bigger than the PDH board is designed for, so there is now a 10dB attenuator between the reflection PD DC block and the RF input to the mixer. 

Here are screenshots of the Inmon channel (which has a gain of ~20) showing a sweep through some PDH signal, and the error signal while in green lock. Huge 60Hz harmonics are still observed. 

TEK00002.PNGTEK00003.PNG

 


Regarding these 60Hz issues, we need to make sure that we remove all situations where long BNCs are chained together with barrel connectors, or Ts are touching other ones. We also should glue or affix the pomona summing box to the shelf, so that its not just laying on the floor.

The concrete next step is to go fiddle with things, and see if we can get the 60Hz noise to go away, then measure the PDH loop and noises again. Hopefully, this should make the ALS much more reliable. 

  10417   Wed Aug 20 21:09:16 2014 ericqUpdateGreen LockingXarm Green PDH

I remeasured all of the noise spectra again today, making sure the input attenuation was as low as it could safely be. I also got a snap of the y green PDH signal; it's fairly larger than I saw the other day, which is good. I used this to calibrate the error signal voltage spectra. 

scopeSweep.jpg

Here are the noise traces for each arm. During these measurements GTRX was about .6, GTRY about 1.0 The Yarm noise doesn't look so good: the error signal is just barely above the mixer+lowpass output noise, and the RMS is plauged by 60Hz lines. (Is this related to what we see in IR TRY sometimes?)

Xspectra.pdfYspectra.pdf

Here are the arms error signals compared directly:

XYcomp.pdf

  10416   Wed Aug 20 18:05:18 2014 JenneUpdateGreen LockingYarm Green PDH - requirement

 

 I calibrated the control signal from Volts to Hz using the rough PZT calibration of 5MHz/V for the Yend NPRO.  

For the error signal, Q said that the Yarm PDH peak-to-peak height was about a factor of 100 smaller than the Xarm, so I used a calibration of 1.9e7 Hz / V.

Then, from Q's Mist simulation including the high Xarm loss, and the plot that he posted in the control room, the CARM linewidth looks like it is about 2pm.  This is the number that I have included on today's plot.  Note though that yesterday I was using a linewidth of about 30pm, which I got from an Optical simulation about a year ago.  I do not know why these numbers come out an order of magnitude different!      The CARM linewidth is actually about 20 pm.  Both Q and I failed at reading log-x plots yesterday.  I have corrected this, and replotted.

Anyhow, here's the Yarm noise spectra calibrated plot:

YPDH_noise.pdf

I have emailed Kiwamu, but haven't heard back from him yet on what the original design considerations were, if he remembered us ever using a boost, etc.  What this looks like to me is that we need to do some serious work to get the noise down.  Maybe fixing the gain peaking and triggering the boost will get us most of the way there?

  10415   Wed Aug 20 16:10:43 2014 ericqUpdateGreen LockingXarm Green PDH

A MIST simulation tells me that the green pdh horn-to-horn displacement is about 1.2nm, or ~18kHz. I used this, along with the scope trace attached to the previous post, to calibrate the mixer output at 193419 Hz per V. (EDIT: I was a little too hasty here. What I'm really after is the slope of the zero crossing, which turns out to be almost exactly twice my earlier naïve estimate. See later post for correct spectra)

For the control signal, I assumed a flat Innolight PZT PM response of 1MHz/V. ( Under 10kHz, it is indeed flat, and this is the region where the control signal is above the servo output noise in yesterday's measurements)

Here are all of the same spectra from last night, with the above calibrations. 

XspectraCombined.pdf

 

Going off Jenne's earlier plot, it looks like the in-loop error signal RMS is ten times bigger than the CARM linewidth. 

  10414   Wed Aug 20 15:31:27 2014 ericqUpdateCOCArm Loss Investigations Continue

 [ericq, Gabriele] 

Summary: After today's meeting, Gabriele and I looked into the arm loss situation, to see if we should really believe the losses that had been suggested by my previous measurements. We made some observations that we're not sure how to explain, and we're thinking about other ways to try and estimate the losses to corroborate previous findings. 


We first looked to see if the ASS had some effective offset, leaving the alignment not quite right. Once ASS'd, we twiddled each arm cavity mirror in pitch and yaw to see if we could achieve higher transmission. We could not, so this suggested that ASS works properly. 

We then looked at potential offsets in the Xarm loop. We found that an input offset of 25 counts increased the transmission, but only very slightly. With this offset adjusted, we confirmed the qualitative observation that locking/unlocking the xarm causes a much bigger change in ASDC than doing the same with the harm.

However, we noted that the ASDC data (which is the DC value of the AS55 RFPD) was quite noisy, hovering around 50 counts. Looking at the c1lsc model, we found that we were looking at direct ADC counts, so the signal conditioning was not so great. We went to the LSC rack and stole the SR560 that had been hooked up as a REFLDC offsetter, and used it to give ASDC a gain of 100, and a LP at 100Hz, since we only care about DC values. We then undid the gain in the input FM; and this calmed the trace down a fair bit. The effects due to each arm locking/unlocking was still consistent with previous observations. 


At this point, we looked at the arm transmission and ASDC signals simultaneously. Normally, when misaligning a cavity, one would expect the reflected power to rise and the transmission to fall.

However, we saw that when misalignment the Yarm in yaw in either direction, or the Xarm in one direction, both the IR transmission and ASDC would fall. This initially made us think of clipping effects. 

So, we checked out the AS beam situation on the AP table. On a card, the beam looks round as we could tell, and the beam spot on AS55 was nice and small. (We tweaked its steering a little bit in pitch to put it at the center of the "falling-off" points) The reflection and transmission falling effect remained. 


At this point, we're not really sure what could be causing this effect. After the reflected beams recombine at the BS, the output path is common, so it's strange that this odd effect would be the same for both arms. 

Lastly, we discussed other ways that we may be able to see if the Xarm really has ~500ppm loss. Since its transmission is ~1.4%, Gabriele estimated that we may be able to see a ~300Hz difference in the arm cavity pole frequency between the two arms, based on the modification of the cavity finesse due to loss. Since we don't currently have the AOM set up to inject intensity noise, we talked about using frequency noise injection to measure the arm cavity poles, though this would be coupled with the IMC pole, but this could hopefully be accounted for.

  10413   Wed Aug 20 04:09:21 2014 ericqUpdateGreen LockingXarm Green PDH

I've made a whole bunch of measurements on the Xarm green situation.

TL;DRs:

  • GTRX was around 0.55 for all of the measurements tonight. 
  • Based on where I saw gain peaking in the CLG, it looked like UGF was 1-2kHz. I cranked the gain to 10kHz, ~20dB gain peaking followed, making it hard to measure. Currently sitting at 5kHz-ish. 
  • Measured CLG with AG4395A, calibrated for injection point response, inferred OLG. 
  • Took various PSDs, still need to calibrate into physically meaningful units. 

Reasonable amounts of time were spent bending the AG4395 to my will; i.e. figuring out the calibration things Jenne and Rana did, finding the right excitation amplitude and profile that would leave the light steadily locked, and finding the right GPIB incantation for getting spectra in PSD units instead of power units. I'm nearing completion of a newer version of AG4395 scripts that have proper units, and pseudo-log spectra (i.e. logarithmically spaced linear sweeps)


Transfer functions

Here is too many traces on one plot showing parts of the OLTF for the x green PDH. One notable omission is the PD response (note to self:check model and bandwidth). The servo oddly seems to have a notch around 100k. My calibration for the CLG injection may not have been perfect, instead of flattening out at 0dB, I had 2dB residual. I tried to correct for it after the fact, assuming that certain regions were truly flat at 0dB, but I want to revisit it to be thorough. I found some old measurements of the Innolight PZT PM response, which claims to be in rad/V, and have included that on the plot. 

Xbodes.pdf

In the end, the mixer and PZT response make it look like getting over 10kHz bandwidth may be tough. Even finding a good higher modulation frequency to be able to scoot the LP up would leave us with the sharp slope in the PZT phase loss, and could cause bad gain peaking. Maybe it's worth thinking about a faster way of modulating the green light?


Noise Spectra

Tomorrow morning, I'll calibrate all the noise spectra I have into real units. These include:

  • In loop error signal and control signal spectra
  • Mixer output spectrum when PD is dark, and when mixer input is terminated
  • Servo out spectrum when PD is dark, and when servo input is terminated

However, looking at the floors, it occurs to me that I may have left the attenuation on the input too high, in an effort to protect the input the PDH box, which rails all the time when not locked to a 00 mode, sometimes even with the input terminated or open. It's kind of a pain that the agilent makes it really hard to see the data when you're in V/rtHz mode, because I should've caught this while measuring :/

I used a scope to capture a pdh signal happening, which will let me transform the mixer output into cavity motion. The control signal goes to the innolight PZT with a ~1MHz/V factor. Here are the uncalibrated plots, for now. 

scopeSweep.jpg

Xspectra.pdf

 

  10412   Wed Aug 20 02:38:41 2014 JenneUpdateGreen LockingYarm Green PDH - requirement

Quote:

* (JCD) Think about this box's purpose in life.  What kind of gain do we need?  Do we need more / less than we're currently getting? NPRO freq noise is 1/f and is 10kHz/rtHz at 1Hz (this is from a plot of an iLIGO NPRO from Rana's thesis, but it's probably similar). Talk to Kiwamu; the noise budget in the paper seems to indicate that we had some kind of boost on or something.  Also, if we need much more gain than we already have, we'll definitely need a different box, maybe the PDH2 box that they have over in WBridge.

It's not so impressive yet, but here's a plot that shows (a) Rana's guess for laser frequency noise, (b) The inferred in-loop version of that noise, (c) The CARM linewidth FWHM, translated to Hz.

For (b), I take the loop that Rana and I measured last night, and I assumed that it continued on forever as 1/f toward low frequency.  Then I do 1/(1+G) to get the closed loop version of the loop (which is a measurement with an artificial line tacked on the end), and multiply this with the laser freq noise, which is also totally artificial.

For (c), I do df/f = dL/L, with f = c/lambda_green, since the rest of the plot is meant to be in green frequency units.

This is my beginnings of trying to come up with a requirement for our green PDH boxes.  We weren't very clear in the MultiColor paper about the nitty-gritty details (obviously), but then Kiwamu didn't expand on those details in his thesis either.  He talks a lot more about the design considerations for the digital ALS loop, which isn't what I want today.  I will send him an email to see if he had any notes that didn't make it into his thesis.

NoiseConsideration.pdf

  10411   Tue Aug 19 23:11:15 2014 JenneUpdateGreen LockingYarm Green PDH

 

 Here is a plot of last night's data with both the control and the error point on the same plot, in Volts.  Q is still working, so I don't have a calibration number yet to get these to Hz.

Note in the control spectrum that we have very significant 60Hz lines.  

ErrAndCtrlSpectra_VoltsPerRtHz.png

EDIT:  I also added a new branch to the DCC Document Tree, and 2 leafs (one for each end).  Here's the ALS PDH servo branch: E1400350

  10410   Tue Aug 19 21:40:44 2014 AndresUpdateIMCNew Optical Setup for the IMC

IMC Calculation and Setup

I have been working in the calculation for improving the Gouy Phase separation between the WFSs. I tried different possible setup, but the three big constrains in choosing a good optical table setup are to have a Waist size that range from 1mm-2mm, the Gouy Phase  between the WFSs have to be greater than 75 degrees and there has to be a steering mirror before each WFS. I will be showing the best calculation because that calculation complies with Rana request of having both WFSs facing west and having the shortest beam path. I approximate the distances by measuring with a tape the distance where the current optics are located and by looking at the picture that I took I approximated the distance where the lenses will be placed. I'm using a la mode for calculating the gouy phase different. I attached a picture of the current optical table setup that we have. Using a la mode, I found that the current gouy phase that we have is 49.6750 degrees.

Now, for the new setup, a run a la mode and found a Gouy phase of 89.3728 degrees. I have to create a two independent beam path: one for the WFS1 and another one for WFS2. The reason for this is that a la mode place everything in one dimension so and since the WFS1 will have a divergence lens in order to increase the waist size, and since that lens should not be interacting with the waist size in the WFS2. We need two beam path for each WFS.  A la mode give us the following solution:

For the beam path of the WFS1

    label                z (m)           type             parameters        
    -----                  -----              ----             ----------        
    MC1                   0              flat mirror          none:           
    MC3                   0.1753     flat mirror          none:           
    MC2                   13.4587   curved mirror    ROC: 17.8700 (m)     
    Lens1                 29.3705   lens                  focalLength: 1.0201 (m)
    BS2                    29.9475   flat mirror          none:           
    First Mirror         30.0237   flat mirror          none:           
    Lens3                30.2000    lens                  focalLength: -0.100 (m)
    WFS1                30.4809    flat mirror         none: 

For the beam path of the WFS2

    label                   z (m)             type             parameters        
    -----                    -----                 ----             ----------        
    MC1                    0               flat mirror          none:           
    MC3                    0.1753      flat mirror          none:           
    MC2                    13.4587    curved mirror    ROC: 17.8700 (m)     
    Lens1                  29.3705    lens                   focalLength: 1.0201 (m)
    BS2                     29.9475    flat mirror          none:           
    Second Mirror    30.2650     flat mirror          none:           
    Lens2                 30.4809     lens                  focalLength: -0.075 (m)
    Third Mirror        30.5698     flat mirror          none:           
    WFS2                30.6968      flat mirror          none:  

I attached bellow how the new setup should look like in the second picture and also I include and attachment of the a la mode code.

 I used Mist to be able to see the read out that we get in the WFSs that take the Mode Cleaner Reflection and the QPD that take the transmitted from MC2. In the following, plots I'm misaligned the each mirrors: MC1, MC2 and MC3. The misalignment are in Yaw and Pitch. I'm dividing the WFSs reading by the total power reflect power, and I'm dividing the QPD for the MC2 transmission by the total transmitted power. In my Mist model, I have a laser of 1W and my EOM is modulated at 30MHz instead of 29.5MHz and the modulation depth was calculating by measuring the applied voltage using and Spectrum analyzer. I using Kiwamu measurement of modulation depth efficiency vs the applied voltage, https://dcc.ligo.org/DocDB/0010/G1000297/001/G1000297-v1.pdf,  I got a modulation depth of 0.6 mrad. I put this modulation depth and I got the following plots: The fourth and fifth attachment are for the current optical setup that we have. The sixth and seventh attachment is for the new optical setup. The eighth attachment is showing the mode cleaner cavity resonating. The last attachment contains the plots of WFS1 vs WFS2, MC2_QPD vs WFS1, MC2_QPD vs WFS3 for each mirror misaligned. The last two attachment are the MIST code for the calculation.

We have all the lenses that we need. I checked it last Friday and if everything is good we will be ready to do the new upgrade this coming Friday. For increasing the power, I check and we have different BS so we can just switch from the current setup the BS. Can you let me know if this setup look good or if I need to chance the setup? I would really love to do this upgrade before I leave.

 

 

 

 

 

 

Attachment 1: ModeCleanerSetup.PNG
ModeCleanerSetup.PNG
Attachment 2: NewOpticalTableSetupForTheModeCleaner.PNG
NewOpticalTableSetupForTheModeCleaner.PNG
Attachment 3: ReduceWFSPathWorkingOn.m.zip
Attachment 4: MIST_WFSsAndQPDReadingForYaw.png
MIST_WFSsAndQPDReadingForYaw.png
Attachment 5: MIST_WFSsAndQPDReadingForPitch.png
MIST_WFSsAndQPDReadingForPitch.png
Attachment 6: MIST_WFSsAndQPDReadingForYawNewSetup.png
MIST_WFSsAndQPDReadingForYawNewSetup.png
Attachment 7: MIST_WFSsAndQPDReadingForPitchNewSetup.png
MIST_WFSsAndQPDReadingForPitchNewSetup.png
Attachment 8: MISTResonanceCavityReflectionAndTransmissionNewSetup.png
MISTResonanceCavityReflectionAndTransmissionNewSetup.png
Attachment 9: 2Dplots.zip
Attachment 10: ModeCleanerCurrentOpticalTableMIST.zip
Attachment 11: ModeCleanerNewSetupMIST.zip
  10409   Tue Aug 19 18:32:40 2014 ericqUpdateGreen LockingYarm Green PDH

Heading to dinner, going to come back for more green fun, but here's a quick update:

Xarm Peak-to-Peak of the PDH signal in the mixer output is about 70mV when GTRX was about 0.4. The sideband-generating function generator has an output of 2V (forgot to note rms or pp)

Yarm Peak-to-Peak of the PDH signal in the mixer output is about 640uV when GTRX was about 0.71. The sideband-generating function generator has an output of 0.091V (forgot to note rms or pp)

The Yarm signal thus correspondingly has a waaay noisier trace. I would've had scope plots to show here, but the scope freaked out about how large my USB drive capacity was and refused to talk to it >:|

This suggests to me that our modulation depth for the Yarm may be much too small, and may be part of our problems with it. 

  10408   Tue Aug 19 01:01:36 2014 Jenne, RanaUpdateGreen LockingYarm Green PDH

[ Rana, Jenne]

We remeasured the Yend PDH box.

When we first started, the green couldn't hold lock to the arm - it kept flickering between modes.  Changing the gain of the PDH box (from 7.5 to 6.0) helped.

We measured a calibration, from our injection point to our measurement point.

The concept was that we'd take the mixer output, and put that into an SR560, and put the swept sine injection into the other input port of the '560, and use A-B.  So, for this calibration, we left A unplugged, and just had the RF out of the 4395 going to input B of the '560.  The 600 Ohm output of the '560 went to the error point input on the PDH box (during normal operation the mixer output is connected directly to the error point input).  The SR560 was set to gain of 1, no filtering.  I don't recall if we were using high range or low noise, but we tried both and didn't really see a difference between them.

We had the 4395 take that calibration out, and then we measured the closed loop gain up to 1 MHz. (Same measurement setup as above, but we connected the mixer out to the input of the SR560 to close the loop, and made sure we were locked on a TEM00 green mode.) Rana used an ipython notebook to infer the open loop gain from our measurement.  Our conclusion is that we don't have nearly enough gain margin in our loop.  We found the PDH box gain knob at 7.5, and we turned it down to 6.0, but the loop is still pretty borderline. We used the high impedance active probe to measure the error point monitor, since we aren't sure that that point can drive a 50 Ohm load.

YPDH_OLG.pdf

We also measured the error point spectra and the control point spectra.  Unfortunately, the saved data from the analyzer (no matter what is on the screen) comes out in spectrum, not spectral density.  So, we need to check our conversion, but right now to get from Watts power to Volts, we do sqrt(50 ohm * data).  We then need to get to spectral density, and right now we're just dividing by the square root of the bandwith that is reported in the .par file. This last step is the one we want to especially check, by perhaps putting some known amount of noise (from an SR785?) into the 4395, and checking that our calibration math returns the expected noise spectrum.

What still needs to be done is to calibrate this into Hz/rtHz.  To do this, we were thinking that we should look at the error point on a 'scope while the cavity is flashing.

Anyhow, here is the uncalibrated error point spectrum.  Purple is a measurement up to 30kHz, with 30Hz bandwidth.  Blue is a measurement up to 300kHz with 300Hz bandwidth.  The gain peaking schmutz above 10kHz sucks, and we'd like to get rid of it.  We also see the same peak at ~150kHz that Q saw earlier today.  We were using the high impedance probe here too.

YPDH_noise.pdf

 We have the data for the control point (all the data files are in /users/jenne/ALS/PDHloops/Yend_18Aug2014), but we haven't plotted it yet.

Things that need doing:

* (JCD) Think about this box's purpose in life.  What kind of gain do we need?  Do we need more / less than we're currently getting? NPRO freq noise is 1/f and is 10kHz/rtHz at 1Hz (this is from a plot of an iLIGO NPRO from Rana's thesis, but it's probably similar). Talk to Kiwamu; the noise budget in the paper seems to indicate that we had some kind of boost on or something.  Also, if we need much more gain than we already have, we'll definitely need a different box, maybe the PDH2 box that they have over in WBridge.

* (EQ, priority 1) Measure and calibrate error point noise down to lower freq for both arms.  What could we win by putting in a boost? If the residual noise is high, maybe the laser isn't good at following arm, so beatnote isn't good length info for the arm, and we can't succeed.

* (EQ, priority 2) Measure TF of PDH box, and a separate measurement of the Pomona box that is between the mixer and the error point - is that eating a bunch of phase?  It's already an LC circuit which is good, but do we really want a 120kHz lowpass when our modulation frequency is roughly 200kHz?  Ask ChrisW - he worked on one of these with Dmass.

* (EQ, priority 2ish) Measure TF of Xend PDH loop (unless you already have one, up to ~1MHz).

* (JCD) Make DCC tree leaf for PDH box #17.  Take photos of box.

  10407   Mon Aug 18 18:33:57 2014 ericqUpdateGreen LockingYarm Green PDH

So far today, I've been working with the Y-end green PDH locking. Using a SR560 to roll off the AG4395A output to take a loop measurement at the servo output, I measured the following OLG, and inferred the CLG from it. The SR560 really helped it getting good coherence without introducing a big offset that changes the optical gain, thus distorting the loop shape, etc. etc. 

yLoop.pdf

You would think this loop looks pretty good, 10k UGF, and 45 degrees of phase margin, gain peaking is sane, and pretty smooth slope. But, the thing still was flipping out of lock while I measured this. 

I suspect shenanigans at >100k. This is motivated by the fact that I've seen some big noise in the error signal around 150k. I don't have a good noise plot right now, because I'm trying to get a scheme going where I stitch together a bunch of 1 decade spectra from the 4395, but the noise floor isn't consistent across each patch (even though the attenuation stays the same, and I confirmed I'm in "noise" mode). I'm working on a loop measurement up there, too, but I haven't been able to get the right filter/amplitude settings yet. 

So, even though this plot is not totally correct (read: wrong and bad), I include it just for the sake of showing the big honking spike of noise at ~150K.  

crap.pdf

 

  10406   Mon Aug 18 09:42:50 2014 KojiSummaryIOOMCREFL PD charcterization

Riju did the measurement of the MCREFL PD.
I found data files in her directory on the control machine.

I was not sure how much was the transimpedance of the DC out.
I assumed the default number from the circuit diagram which was 66.7Ohm.
This may cause the error in absolute caribration of the transimpedance but the shape does not change.

The RF preamp is gain-peeking at 250MHz.

Here is further characterization of the PD response.
As you can see in the second attachment, the 3dB cut off of the resonance is about 2.3MHz.

The game plan file in dropbox was also modified.

Attachment 1: MCREFLPD_transimpedance.pdf
MCREFLPD_transimpedance.pdf
Attachment 2: MCREFLPD_transimpedance_zoom.pdf
MCREFLPD_transimpedance_zoom.pdf
  10405   Fri Aug 15 20:38:17 2014 ericqUpdateGeneralELOG dump

 A few things that I have neglected to ELOG yet:

  • scripts/offsets/LSCoffsets is a new script that uses ezcaservo to set FM offsets of our LSC PDs. It still warns about large changes, and lets you revert. It reads the FM gain to pick the right gain for the ezcaservo call. 

  • MC refl DC was all over the place today, and has recently been "fuzzier" on the wall StripTool than I like. I touched the MC2 pointing a little bit, and the WFS seemed to find a sweet spot where the refl got steady back at around and under 0.5. I then ran "offload WFS" to try and stay there. 

  • Incidentally, the PMC transmission drifted up to 0.81 at some point today. This is weird, since not too long ago, we were not able to reach this level even with careful alignment. This coincided with the MC power being back up to ~17k, and arms locking at around 0.95. 

  • Last week I quickly tried cranking up the x-end green modulation frequency to ~1.3MHz (corresponding to a notch in the PZT AM response), and using a 550k lowpass on the mixer output, instead of a 70k, to try to buy more phase and increase the UGF. It didn't work. I didn't have a way to tune the mixer phase angle, and the mixer output was super noisy, but there were instants where I could convince myself that a mode was briefly locked to the arm... I'm going to do the Right Thing and characterize the loop properly, to figure out how to get at least 10kHz of control bandwidth out of these things. 

  10404   Fri Aug 15 20:26:37 2014 ericqUpdateGeneralGame plan

Quote:

Q already did the tweak up of the PSL SHG crystal alignment.  HE SHOULD ELOG ABOUT THIS.  What was the final power of green that you got?  Do we have any record of a previous measurement to compare to?

As Jenne mentioned, I did this. 

Specifically, I first tweaked the mirror pointing the IR into the SGH in pitch and yaw to maximize the green power, and then adjusted the little set screws on the side of the SHG to maximize further. Power after the harmonic separator was of order 150uW. On the Y Green BBPD, I got ~48uW, instead of the 40uW Rana, Jenne, and myself saw the other night. 

HOWEVER,

now that I look through old ELOGs, I find some posts by Kiwamu saying the power should be around 650uWand that he was able to get 640uW out. So: I should do this again, systematically, more carefully, etc., etc. (Linked ELOG also states that optimum SHG temperature is alignment dependent...)

 

 

  10403   Fri Aug 15 17:24:44 2014 HarryUpdateGeneralFiber Temp.

 Earlier today Q and I somewhat resurrected my old PER measurement setup so I could run the temperature characterization experiment.

Unfortunately, when I tried to use the fiber illuminator, no light came from the other end, causing me to fail my primary goal for the summer of "don't break anything." The fiber has been re-spooled and labeled appropriately. Also sorry.

In addition to this, Q and I scavenged parts from the telescopes on the PSL and Y End tables, which were either not functional, or needed to have their mode matching adjusted, since we're using the non-PM fibers for FOL, which have a different numerical aperture, and thus slightly different output modes.

Specifically, this is involved removing the rotational mounts, and appropriate beam dumping.

My "calorimeter" still remains intact, in case anyone wants to make this measurement in the future, as this is my last day in the lab.

It's also effective at keeping drinks cold, if you'd rather use it for that.

  10402   Fri Aug 15 14:35:57 2014 ericqUpdateLSCTRY mystery offset gone

One question answered, but another raised. The offset came from LSC-TRY switching to the ETMY-QPD signal from ETMY-TRY (Hi gain pd). 

BUT WHY

TRYmystery.png

  10401   Fri Aug 15 14:09:21 2014 JenneUpdateLSCTRY mystery offset gone

Again unknown, but about 6 hours ago (so ~8am) the offset disappeared. 

Here's a 1-day trend:

TRY_0pt3_offset_gone.pdf

  10400   Fri Aug 15 13:29:31 2014 JenneUpdateGeneralGame plan: 15 Aug

40mToDoList.pdf

The game plan graffle file is now in the 40m dropbox, so anyone can edit it.  Please just make sure to keep the date in the top right corner accurate.

  10399   Fri Aug 15 02:05:55 2014 JenneUpdateLSCALS in-loop spectra

 Not sure why, but Rana and I didn't see the super high Xarm noise with ALS that we reported last night (elog 10382).  

The in-loop ALS noise seems fine.  The out of loop measurement while the ALS is locked is a little tricky, since ALS hold the arms within the POX/POY linear ranges.  

Here is the in-loop noise:

ALS_XY_inloop_noise_14Aug2014_lowBeatFreq.pdf

  10398   Fri Aug 15 01:31:44 2014 JenneUpdateASCPOP QPD to ETMs model wiring complete

I have added a few things to the ASS model, and the ASC sub-block, so that we can send POP QPD information down to the ETMs for CARM angular control after we've reduced the CARM offset and gotten some carrier buildup.  I did not remove our ability to actuate on PRM, so that we can still play with it in PRMIsb cases.

The input matrix has been expanded so that it can send signals to new CARM_YAW and CARM_PIT filter banks.  The corresponding filter banks have been created.  The output matrix was also expanded to take in the 2 new servo outputs, and so it can send signals to both ETMs, pitch and yaw.  I did not include any triggering logic for this new CARM situation, since I assume we'll just turn it on and off with our scripts.  (We haven't really been using the triggering capability of the PRM ASC either lately, although it's all still there).  I added the inputs and outputs of the CARM servos to the list of acquired channels.

The ASC sub-block:

ASC_newCARM.png

I also modified the top level of the ASS model.  This was just a simple addition of summing nodes for the ETMs, similar to what was already in place for the PRM, so that we can send both the ASS dither alignment signals and the ASC servo control signals to the optics.

The ASS top level:

ASSoverview_newCARM.png

I also quickly modified the ASC screen to expose all of the new options:

ASC_screen.png

The ASS model was compiled, and restarted.  As usual, this temporarily removes the biases on the input pointing tip tilts, but the pointing seems to have come back without any trouble.

  10397   Thu Aug 14 23:19:49 2014 ranaSummaryLSCETM Violin fundamental filters moved to LSC

 We used to do violin mode and test mass body mode notches in the SUS-LSC filter modules. Now we want them balanced in the LSC and triggered by the LSC, so they're in the filter modules which go from the the LSC output matrix to the SUS.

01.png

Today, we were getting ETM violin mode ringups while doing ALS hunt and so we moved the bandstops into the LSC. I also changed the bandstop from a wide one which missed the ETMX mode to a double bandstop which gets both the ETMX and the ETMY mode. See attached image of the Bode mag.

03.png

  10396   Thu Aug 14 22:58:59 2014 rana, jenneSummaryGreen LockingALS DIFF tuning

 We've been having trouble tuning the ALS DIFF matrix. Trying to see if the MC2 EXC can be cancelled in ALS DARM by adjusting the relative gains in ALSX and ALSY Phase Tracker outputs.

There's a bunch of intermittent behavior. Between different ALS locks, we get more or less cancellation. We were checking this by driving MC2 at ~100-400 Hz and checking the ALS response (with the ALS loops closed). We noticed that the X and Y readbacks were different by ~5-10 degrees and that we could not cancel this MC2 signal in DARM by more than a factor of 4-5 or so. In the middle of this, we had one lock loss and it came back up with 100x cancellation?

Attached is a PDF showing a swept sine measurement of the ALSX, ALSY, and DARM signals. You can see that there is some phase shift between the two repsonses leading to imperfect cancellation. Any ideas? Whitening filters? HOM resonance? Alignment?

Attachment 1: sweep.pdf
sweep.pdf sweep.pdf
  10395   Thu Aug 14 22:31:12 2014 JenneUpdateLSCTRY gets mystery offset

I don't know why, but TRY has somehow gotten a 0.3 count offset in the last hour. 

Rana and I are witnesses for each other that neither of us has gone into the IFO room in the last several hours (and we're the only ones here).  For some reason though, the TRY PD now has a 0.3 count offset.  We have been doing some ALS locks, but we have not run the offset script in the last several hours.  Closing the green shutter doesn't change things, and we still see the offset when the MC loses lock, so it's not to do with the end or the PSL laser.  We haven't been in there, so there hasn't been a change in the room lights. 

TRY_0pt3_offset.pdf

  10394   Thu Aug 14 22:16:02 2014 JenneUpdateSUSViolin Mode filters for ETMs

The instigator of this was that we were seeing ring-ups of ETMs during our ALS locks this evening.  We measured the ETMY violin resonance to be 624.10 Hz, and Rana found an elog saying that the ETMX was around 631 Hz, so we made a 2 notch filter and added it to FM4 of the LSC-SUS filter banks for both ETMs. 

For the ETMY resonance, we measured the frequency in the DARM spectrum, and when we looked at the FINE_PHASE_OUT channels, the resonance was only in the Yarm sensor.  So, we conclude that it is coming from ETMY.

Also in the realm of filter modules, the FM3 boost for CARM, DARM, XARM and YARM was changed from zero crossing to ramp with a 1sec ramp time.

  10393   Thu Aug 14 20:52:36 2014 ranaUpdateWikiViolin Mode table added to Wiki

Mech Resonance Wiki

I've updated the wiki by trawling the elog for violin entries. Please keep it up to date so that we can make violin notches.

 

  10392   Thu Aug 14 19:33:00 2014 JenneConfigurationIOOMoved MC2 spot

Last night, and again just now, I used the ./MC2_spot_[direction] scripts to center the MC2 spot on the trans QPD.  The MCWFS handled overall alignment to correct for the fact that the ratios in the script aren't perfect.  When I was finished, I ran the MC WFS relief script from the WFS screen.  Last night, and again today, things had drifted until the yaw spot was more than 0.5 counts off.

  10391   Thu Aug 14 19:23:25 2014 ranaHowToIOOHow do I set the FSS offset to make the PZT voltage start at the right place?

 When the IMC locks, we want the FAST OUT of the TTFSS box to be close to zero volts. We also want the control signal from the MC Servo board to be close to 0 V. How to set this up?

With the IMC locked, we just servo the FSS input offset to minimize the MC board output :

ezcaservo -r C1:IOO-MC_FAST_MON -g 0.1 -t 10 C1:PSL-FSS_INOFFSET

I would have used "CDSUTILS", but that seems to have some sort of ridiculous bug where we can't have prefixes on channel names, even on the command line. 

  10390   Thu Aug 14 18:31:45 2014 ericqUpdateLSCLSC Modeling Update

 Based on the game plan, I have created a slew of updated pretty plots about our signals and loops. 

First: With measured arm losses, when do we start to see REFL DC dip? At what arm buildup powers? 

I updated my MIST model with the arm losses I've measured (Y:130ppm, X:530ppm), and some measured transmissions from the wiki, vs. the design parameters, as I used to have. Here is the DC sweep plot which is now hanging up in the control room. 

dcSweep.pdf

In this plot, I also calculated what MIST thinks the full arm power buildup will be as compared to our single arm locking, and I get something of order 200, rather than the 600 we've tossed around in discussions. Nothing else is very different in this plot from the old version; though the REFLDC dip is a little bit wider. 

Now, here are some radiation-pressure inclusive sensing transfer functions, for the anti-spring case (which in Rob's day was easier to lock for unknown reasons):

carm2TRX.pdfcarm2REFLDC.pdf

carm2REFL11.pdfdarm2AS55Q.pdf


Next: Include new AO path TFs into CM model Look at possibilities for engaging AO path 

With these TFs, and the recently measured+fit new AO TF, here are the open loop gains of the slow, digital, SqrtInv-sensed MCL CARM and fast, analog, REFLDC-sensed AO CARM loops for the region of offsets we've achieved and a little lower. The slow digital loop includes the 1k LP that we have used in the past, in addition to the normal CARM filters. I still need to figure out the right sequence of ( offset reduction / crossover frequency motion / overall gain adjustment ) that gets the coupled cavity resonance solidly within the loop bandwidth. 
 
MCLcarmLoop.pdfAOcarmLoop.pdf

 

  10389   Thu Aug 14 18:10:46 2014 HarryUpdateGeneralFiber Temperature Effects Setup

Purpose

We want to characterize the sort of response the fibers have to temperature gradients along them (potentially altering indices of refraction, etc.)

Experimental Setup

I have constructed a sort of two chambered "calorimeter" (by which I mean some coolers and other assorted pieces of recycling.)

The idea is that half of the length of PM fiber resides in one chamber, and the other in the other.

One chamber will remain at an uncontrolled, stable temperature (as measured by thermocouple probe) while the other's temperature is varied using a heat gun.

Using this setup, one can measure losses in power, and effects on polarization within the fiber.

Caveat

This is currently living on the electronics bench until tomorrow morning, and is a little fragile, just in case it needs to be moved.

Attachment 1: tempAffectsSetup.zip
  10388   Thu Aug 14 18:05:05 2014 JenneUpdateGeneralUpdated game plan

Quote:

- Oplevs for PR2, PR3 => Almost impossible.

 Because of the limited table space inside?  That's the main reason I can think of that this method is hard.  Am I missing something?

  10387   Thu Aug 14 18:02:11 2014 KojiUpdateGeneralUpdated game plan

Got the idea of ASC.

- Oplevs for PR2, PR3 => PR2 seems OK. PR3 almost impossible. well turned out not too crazy. We need outside electronics.

- RF QPD => not trivial and very technical but possible. All outside work.

- Better TT => might be a good solution.

  10386   Thu Aug 14 15:51:37 2014 JenneUpdateGeneralUpdated game plan

Quote:

 - ALS

End PDH UGF improvement / post mixer LPF investigation (with in 2 weeks)

- MC/FSS

Riju measured the MC REFL PD transimpedance. See ELOG and related.

- ASC

Why do we want to see less PRM motion? I thought PRC motion was causing
LSC issue of the central part. We wanted to maximize the PRM effect, don't we?
(Or is this to supress ETM motion during full lock?)

 End PDH - good point, thanks.

ASC - Yes, this is so that we can use the POP QPD to feed back to the common ETMs after the CARM offset is already quite small.  We will not use POP DC QPD for PRC any more. 

Also, for future PRC ASC, I keep coming back to this in my head, but maybe it is less painful to install oplevs for PR2, PR3 than it would be to make an RF QPD.  Neither is going to be trivially easy.  But if we had sensors of the tip tilt motions, we could feed all of that back to the PRM to stabilize the PRC.

  10385   Thu Aug 14 15:42:29 2014 KojiUpdateGeneralUpdated game plan

 - ALS

End PDH UGF improvement / post mixer LPF investigation (with in 2 weeks)

- MC/FSS

Riju measured the MC REFL PD transimpedance. See ELOG and related.

- ASC

Why do we want to see less PRM motion? I thought PRC motion was causing
LSC issue of the central part. We wanted to maximize the PRM effect, don't we?
(Or is this to supress ETM motion during full lock?)

  10384   Thu Aug 14 15:10:47 2014 AndresUpdateIMCCalculation for the input mode cleaner

Quote:

Can you please give us some more details on how this design was decided upon? What were the design considerations?

It would be nice to have a shorter path length for WFS2. What is the desired spot size on the WFS? How sensitive are they going to be to IMC input alignment? Are we still going to be recentering the WFS all the time?

 I did the calculation, and I reduced the beam Path. In my calculation, I restricted the waist size at the WFSs to be between 1mm-2mm also the other parameter is that the Gouy Phase different between the WFSs have to be 90 degrees. I also try to minimize the amount of mirrors used. I found the Gouy phase to be 89.0622 degrees between the WFSs and the following table shows the solution that I got from a la mode:

 

  label                         z (m)                   type               parameters         
    -----                         -----                    ----                  ----------         
    MC1                    0                        flat mirror           none:            
    MC3                    0.1753               flat mirror           none:            
    MC2                   13.4587              curved mirror    ROC: 17.8700 (m)       
    Lens1                 28.8172              lens                   focalLength: 1.7183(m)
    BS2                    29.9475              flat mirror           none:            
    First Mirror         30.0237              flat mirror           none:            
    Lens3                 30.1253              lens                  focalLength: -0.100 (m)
    Lens2                 30.1635              lens                 focalLength: 0.1250(m)
    WFS1                 30.2269              flat mirror         none:            
    Second Mirror    30.2650              flat mirror         none:            
    Third Mirror       30.5698              flat mirror         none:            
    Lens4                30.8113              lens                  focalLength: -0.075 (m)
    WFS2                31.0778              flat mirror         none:     
       

In the first image attached below is the a la mode solution that show the waist size in the first WFS, and I used that solution to calculate the solution of the waist size for the second WFS, which is shown in figure 2. I photoshop a picture to illustrate how the new setup it supposed to look like. 

Attachment 1: SolutionForTheModeCleanerSetup00.png
SolutionForTheModeCleanerSetup00.png
Attachment 2: SolutionForTheModeCleanerSetup11.png
SolutionForTheModeCleanerSetup11.png
Attachment 3: PossibleSetupForModeCleaner.PNG
PossibleSetupForModeCleaner.PNG
Attachment 4: alaModeSolution.zip
  10383   Thu Aug 14 14:58:03 2014 JenneUpdateGeneralUpdated game plan

2014_Aug_14.pdf

(Updated as of 4pm)

  10382   Thu Aug 14 02:51:46 2014 JenneUpdateGeneralGame plan

[Jenne, Rana]

* Decided that earlier mode hop scan won't give us the information that we were hoping for.  We need to think about where we can actually see the frequency change.  Can we use the IR beatnote that we will soon have to do this?  We'd only be able to scan one laser temp at a time, but that's okay.  Leave, say, the PSL temperature alone, and scan one of the end laser temps.  Using the PSL as the reference, we will be able to see if the frequency of the end laser goes crazy and jumpy as we pass through a certain temp.  Then, repeat while holding the end laser constant and scan the PSL.  Thoughts?

* Meditated on PSL oplev servo, but I need to make a Matlab script that can evaluate different loops according to a cost function based on elog 9690

* Aligned IFO to IR, then greens to arms (got back to 0.9 for GTRY, but only about 0.5 for GTRX, with the PSL green shutter closed).  Then aligned green beams on the PSL table, since the PSL green pointing had changed a bit from Q's crystal alignment tweak-up earlier today.  Beatnotes are nice and big (see elog 10381 - The Yarm is the larger beatnote, and the Xarm is the smaller one.)

* Was not able to lock ALS comm/diff and hold long enough to get both arms to IR resonance.  Also, saw that TRY's RIN was more than 50%(!!!).  We took a look, and there seems to be much more low frequency noise than there was when the spectrum in the control room was taken for the multicolor metrology paper:

Y_ALS_FINE_PHASE_OUT_noisy_13Aug2014.pdf

* Tried to balance the ALS comm/diff input matrix, with not a lot of success.  First of all, it looks like the Xarm has overall about 10 times more noise!  We were exciting MC2 in position (~88 Hz, about 130 counts I think), and then looking at DARM_IN1 for the peak.  When DARM_IN1 was just one of the 2 ALS error signals (i.e. one matrix element set to zero), versus when both matrix elements were set to 1, we saw a factor of only about 3 in reduction of the peak height.  We were hoping to have better cancellation of this pure CARM signal in the DARM channel.  The Xarm green PDH loses lock every ~5 or 10 minutes, and when we relock it, this cancellation seems different, so we want to try again tomorrow when the ALS is locked on comm / diff, rather than just the free running ALS that we have now.  Although, if the balance of the input matrix changes lock-to-lock, we may need to consider redoing the green PSL table layout so we get a pure DARM beatnote signal like they have at the sites.

* We want to change how the watch script for ALS works, although this is a low-priority task.  Rather than looking at the control signal, we should maybe look at the sum of all the coil outputs, multiplied by a pendulum TF, and use that as a rough displacement sensor.  We want to be careful of pushing too hard at low frequencies, but we want to allow higher frequency actuation without having the watch script shut things down.

* Also, I should put on the to-do list the revamp of the ALS find IR resonance script.

  10381   Wed Aug 13 23:58:49 2014 ranaHowToComputer Scripts / ProgramsHP8591E spectrum analyzer remote scan

Quote:

The script for running continuous scans on HP 8591E spectrum analyzer is located at scripts/general/netgpibdata/HP8591E_contdScan.py

There was no such script in the directory when I looked today, but I found one called HP8591E. Of course, it didn't run because it hadn't been tested from the scripts directory and pointed to some /users/nichin/ stuff.

I modified a couple of lines and then committed it and the default .YML parameter file to the SVN. It runs and produces plots continuously from the scripts directory.

*** also, as you can see, we have mostly recovered the green beat amplitudes after yesterday's FLL attack on the ALS ***

Attachment 1: HP8591E_View.pdf
HP8591E_View.pdf
  10380   Wed Aug 13 23:08:17 2014 ranaUpdateIOOFSS box TFs

As EQ pointed out recently, we're injecting into the FSS error point just after an RF pi filter, but before the VGA. We wondered what the weird filter impedance was doing to our signal if we inject after it. I used LISO to model this FSS common section and attach the plots.

The first plot shows the TF between the Test 1 input and the AD602 VGA input. This is NOT the input that we are actually using.

The second plot shows the TF between the IN1 port (which we are actually using) and the VGA input.

Neither of them shows the 1 MHz bump that we see in the measurements, so I suspect that the board has been modified...the hunt continues. We've got to pop the top of the TTFSS and take photos and measure from IN1 to VGA input.

** FSScomm.fil is now in the LISO SVN. The following command line will run it with two different cases and cat the PDF files into one. If you use an auto-refresh PDF viewer like Okular or Mac Preview, its a nicer display than the usual GNUplot window:

./mfil FSScomm.fil; sleep 1; pdftk FSScomm_run*.pdf cat output FSScomm.pdf

Attachment 1: FSScomm.pdf
FSScomm.pdf FSScomm.pdf
  10379   Wed Aug 13 22:01:57 2014 ranaUpdateIMCCalculation for the input mode cleaner

Nic, Andres, and I discussed some more about the MC WFS project today. We want to shorten the proposed WFS2 path. Andres is going to explore moving the 2" diameter lens in coming up with layouts. We also want the WFS to face west so that we can see the diode face with an IR viewer easily and dump the reflected beams in the razor dumps.

We wondered about fixing the power levels and optical gain:

  1. What is the MC modulations depth? What would happen if we increase it a little? Does anyone know how to set it? Will this help the MC frequency noise?
  2. What is the max power on the WFS? I guess it should be set so that the power dissipation of the detector is less than 1 W with the MC unlocked. So P_diss = (100 V)*(I_tot), means that we should have less than 10 mA or ~50 mW when the MC is unlocked.
  3. Another consideration is saturation. The RF signals are tiny, but maybe the DC will saturate if we use any more power. The quadrants are saturated when unlocked and ~200 mV locked. According to D990249, the DC gain in the head is 1000 V/A. The measured power levels going into the heads (w/ MC unlocked) are: P_WFS1 = 4.9 mW and P_WFS2 = 7.7 mW. We don't have control of the DC gain, but there is a 10x and 100x switch available inside the demod board (D980233). From these numbers, I figure that we're in the 100x position and so the effective DC gain between photocurrent and the DC readback voltages is 100 kOhm. Therefore, we are in no danger of optical or electronics saturation. And the unlocked photocurrent of ~40/100000=0.4 mA => 0.04 W heat generated in the diode, so we're OK to increase the power level by another factor of 2-4 if we want.
  4.  We noticed that the ADC inputs are moving by ~50 counts out of 65000, so we're doing a really bad job of signal conditioning. This was previously noticed 6 years ago but we failed to follow up on it. Feh.

While checking this out, I converted the McWFS DC offsets script from csh to bash and committed it to the SVN. We need to remove the prefix 'feature' that Jamie has introduced to cdsutils so that we can use C1 again.

 

  10378   Wed Aug 13 19:23:09 2014 JenneUpdateLSCPSL, Aux laser mode hop check

This afternoon Q helped me put in some temporary PDs for checking for any mode hopping behavior in our 3 main lasers. 

Q helped me install PDA55s on each of the lasers (I did the ends, he did the PSL) so that we could do the mode hop temperature check.  For the Yend, I took the leakage transmission through the first Y1 steering mirror after the laser. This beam was dumped, so I replaced the dump with a PDA55. For the Xend, the equivalent mirrors are too close to the edge of the table, so I put in a spare Y1, and reflect most of the light to a beam dump.  The leakage transmission then goes to a PDA55.  Note that for both of these cases, no alignment of main laser path mirrors was touched, so we should just be able to remove them when we're through.  For the PSL, I believe that Q took the rejected light from one of the PBSes before the PMC. 

The end temporary PDs are using the TRX / TRY cables, so we will be looking at the C1:LSC-TR[x,y] channels for the power of the end lasers.  The PSL's temporary PD is connected to the PMC REFL cable.  For the end PDs, since I had filter banks available, I shuttered the end lasers and removed the dark offset.  I then changed the gains to 1, so the values are in raw counts.  The usual transmission normalization gains are noted in one of the control room notebooks.

I did a slow ezcastep and ramped the temperature of all 3 lasers over about an hour.  Since we usually use the PSL around FSS slow slider value of zero, I swept that from -10 to +10.  Since we usually use the Xend laser at around 10,000 counts, I swept that from 0 to 20,000.  For the Yend laser, it is usually around -10,000 counts, so I swept it from -20,000 to 0.  ezcastep -s 0.2 C1:ALS-X_SLOW_SERVO2_OFFSET +1,20000 C1:ALS-Y_SLOW_SERVO2_OFFSET +1,20000 C1:PSL-FSS_SLOWDC +0.001,20000

I was looking for something kind of similar to what Koji saw when he did this kind of sweep for the old MOPA (elog #2008), but didn't see any power jumps that looked suspicious.

Here is the PSL:

ModeHopCheck_PSL_13Aug2014.pdf

The Xend:

ModeHopCheck_Xend_13Aug2014.pdf

And the Yend:

ModeHopCheck_Yend_13Aug2014.pdf

  10377   Wed Aug 13 17:37:43 2014 JenneUpdateGeneralGame plan

2014_Aug_13.pdf

Here's the game plan for things that we need to do to get this IFO locked up. 

Red is for things that should be done today, or tomorrow if they don't get finished today (eg. laser mode hopping temperature check).  Orange is for things that will become red once the current red things are gone (eg. inferring the POP QPD gouy phase, and moving it to minimized PRM information).  Green is for things that we'd like to do, but aren't high priority (eg. X green mode matching).  Blue is for things that we should remember, but not plan on working on soon (eg. putting PZTs on the Yend table for green).

TODAY so far:

Q already did the tweak up of the PSL SHG crystal alignment.  HE SHOULD ELOG ABOUT THIS.  What was the final power of green that you got?  Do we have any record of a previous measurement to compare to?

Q helped me install PDA55s on each of the lasers (I did the ends, he did the PSL) so that we could do the mode hop temperature check.  For the Yend, I took the leakage transmission through the first Y1 steering mirror after the laser. This beam was dumped, so I replaced the dump with a PDA55. For the Xend, the equivalent mirrors are too close to the edge of the table, so I put in a spare Y1, and reflect most of the light to a beam dump.  The leakage transmission then goes to a PDA55.  Note that for both of these cases, no alignment of main laser path mirrors was touched, so we should just be able to remove them when we're through.  For the PSL, I believe that Q took the rejected light from one of the PBSes before the PMC.  He mentioned that he bumped something, so had to realign the beam into the PMC, but that he was able to get the transmission back up to 0.802, when we were seeing it in the mid 0.7's for the last several days.

The end temporary PDs are using the TRX / TRY cables, so we will be looking at the C1:LSC-TR[x,y] channels for the power of the end lasers.  The PSL's temporary PD is connected to the PMC REFL cable.  For the end PDs, since I had filter banks available, I shuttered the end lasers and removed the dark offset.  I then changed the gains to 1, so the values are in raw counts.  The usual transmission normalization gains are noted in one of the control room notebooks.

I did a slow ezcastep and ramped the temperature of all 3 lasers over about an hour.  I'll write a separate elog about how that went.

  10376   Wed Aug 13 16:12:55 2014 HarryUpdateGeneralFOL Layout Diagram

Per Q's request, I've made up a diagram of the complete FOL layout for general reference.

FOLLayout2.png

  10375   Wed Aug 13 13:08:24 2014 ranaUpdateIMCCalculation for the input mode cleaner

Can you please give us some more details on how this design was decided upon? What were the design considerations?

It would be nice to have a shorter path length for WFS2. What is the desired spot size on the WFS? How sensitive are they going to be to IMC input alignment? Are we still going to be recentering the WFS all the time?

ELOG V3.1.3-